OpenStudy (anonymous):

Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)} A. {(-2, -8), (2, 4), (4, 7)} B. {(-2, -8), (0, -3), (4, 7)} C. {(-2, -8), (0, -3), (1, -1)} D. {(1, -1), (2, 4), (4, 7)}

OpenStudy (anonymous):

Whoa! Let me think on this one..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What does replacement set mean to do?

OpenStudy (anonymous):

not really sure

OpenStudy (anonymous):

Okay! Hold on just one second, I'lll figure it out. What math are you in?

OpenStudy (anonymous):

algebra a

OpenStudy (anonymous):

Oh! Okay! I'll try to remember what to do. Been 3 years since I've done Algebra

OpenStudy (anonymous):

I"m pretty sure what you do for these, are plug it in and solve for both x and y for example the first one is: 5x-2y=6 (-2,-8) -2=x -8=y. I'll show the work

OpenStudy (anonymous):

 then do the same for solving for x